# Nonuniqueness of Leray–Hopf solutions to the unforced incompressible 3D Navier–Stokes Equation

### v1

@@ Appendix G Proof of Propositions 3 and 4
$|\xi|U_{\operatorname{near}}\in L^{2},\quad\quad|\xi|\nabla U_{\operatorname{near}}\in L^{2}.$ (G.13)
Step 2: $H^{1}$-level estimate, that $|\xi|^{2}\nabla U_{\operatorname{near}},|\xi|^{2}\nabla^{2}U_{\operatorname{near}}\in L^{2}$. Take derivatives of (G.11) to get
$\nabla U_{\operatorname{near}}-\frac{1}{2}\xi\cdot\nabla^{2}U_{\operatorname{near}}+(\nabla\widetilde{U})^{2}+\widetilde{U}\cdot\nabla^{2}\widetilde{U}+\nabla^{2}P-\Delta\nabla U_{\operatorname{near}}=\nabla F.$ (G.14)
Now we take inner products of (G.14) with $\chi_{R}\phi_{L}|\xi|^{\alpha}\nabla U_{\operatorname{near}}$, where the cutoff radius $R$ is much larger than $L$, and we get the following estimates as we take $R\to\infty$. We will first take $\alpha=0$ to derive an intermediate result $|\xi|\nabla^{2}U_{\operatorname{near}}\in L^{2}$, and then take $\alpha=2$ to conclude.
@@
When $\alpha=0$, we have the estimate via (G.13) and the decay of $U_{\operatorname{far}}$:
$\||\xi|^{\frac{2+\alpha}{2}}(\nabla\widetilde{U})^{2}\|_{L^{2}}\lesssim\||\xi|^{\frac{1}{2}}\nabla U_{\operatorname{far}}\|_{L^{4}}^{2}+\||\xi|\nabla U_{\operatorname{near}}\|_{L^{2}}\|\nabla U_{\operatorname{near}}\|_{L^{\infty}}\lesssim 1.$
and when $\alpha=2$, we have the estimate by Gagliardo–Nirenberg that
@@
$\||\xi|^{\frac{2+\alpha}{2}}(\nabla\widetilde{U})^{2}\|_{L^{2}}\lesssim 1+\||\xi|^{\alpha/2}\nabla^{2}U_{\operatorname{near}}\|_{L^{2}}^{3/2}$ (G.15)
@@
The pressure term is controlled by (G.12) using Calderón–Zygmund theory and the fact that $|\xi|^{\alpha}$ is an $A_{2}$-weight for $3<\alpha<3$ (see Thm. 5.24 in [24]), so that for $\alpha=0,2$, we have by (G.15) that
@@
Finally the forcing term has the trivial estimate by Cauchy-Schwarz
@@
$|\xi|^{2}\nabla U_{\operatorname{near}}\in L^{2},\quad\quad|\xi|^{2}\nabla^{2}U_{\operatorname{near}}\in L^{2}.$ (G.16)

### v2

@@ Appendix F Proof of Propositions 3 and 4
$|\xi|U_{\operatorname{near}}\in L^{2},\quad\quad|\xi|\nabla U_{\operatorname{near}}\in L^{2}.$ (F.13)
Step 2: $H^{1}$-level estimate, that $|\xi|^{2}\nabla U_{\operatorname{near}},|\xi|^{2}\nabla^{2}U_{\operatorname{near}}\in L^{2}$. Take derivatives of (F.11) to get
$\nabla U_{\operatorname{near}}-\frac{1}{2}\xi\cdot\nabla^{2}U_{\operatorname{near}}+(\nabla\widetilde{U})^{2}+\widetilde{U}\cdot\nabla^{2}\widetilde{U}+\nabla^{2}P-\Delta\nabla U_{\operatorname{near}}=\nabla F.$ (F.14)
Now we take inner products of (F.14) with $\chi_{R}\phi_{L}|\xi|^{\alpha}\nabla U_{\operatorname{near}}$, where the cutoff radius $R$ is much larger than $L$, and we get the following estimates as we take $R\to\infty$. We will first take $\alpha=0$ to derive an intermediate result $|\xi|\nabla^{2}U_{\operatorname{near}}\in L^{2}$, and then take $\alpha=2$ to conclude.
@@
When $\alpha=0$, we have the estimate via (F.13) and the decay of $U_{\operatorname{far}}$:
$\||\xi|^{\frac{2+\alpha}{2}}(\nabla\widetilde{U})^{2}\|_{L^{2}}\lesssim\||\xi|^{\frac{1}{2}}\nabla U_{\operatorname{far}}\|_{L^{4}}^{2}+\||\xi|\nabla U_{\operatorname{near}}\|_{L^{2}}\|\nabla U_{\operatorname{near}}\|_{L^{\infty}}\lesssim 1,$
and when $\alpha=2$, we have the estimate by the Gagliardo–Nirenberg inequality that
@@
$\||\xi|^{\frac{2+\alpha}{2}}(\nabla\widetilde{U})^{2}\|_{L^{2}}\lesssim 1+\||\xi|^{\alpha/2}\nabla^{2}U_{\operatorname{near}}\|_{L^{2}}^{3/2}.$ (F.15)
@@
The pressure term is controlled by (F.12) using the Calderón–Zygmund theory and the fact that $|\xi|^{\alpha}$ is an $A_{2}$-weight for $3<\alpha<3$ (see Thm. 5.24 in [26]), so that for $\alpha=0,2$, we have by (F.15) that
@@
Finally the forcing term has the trivial estimate by the Cauchy-Schwarz inequality
@@
$|\xi|^{2}\nabla U_{\operatorname{near}}\in L^{2},\quad\quad|\xi|^{2}\nabla^{2}U_{\operatorname{near}}\in L^{2}.$ (F.16)